OpenStudy (anonymous):

What is 24x^2y^6 - 16x^6y^2 + 4xy^2 divided by 4xy^2? Can someone please help ?

OpenStudy (whpalmer4):

what is 24x^2y^6 divided by 4xy^2?

OpenStudy (anonymous):

6xy^4

OpenStudy (whpalmer4):

Okay, so that's the first term of the quotient. Now we multiply the divisor (4xy^2) by that, and subtract the result from the dividend (24x^2y^6-16x^6y^2+4xy^2). What do you get when you do that?

OpenStudy (anonymous):

So I divide 24xy^6 from 24x^2y^6-16x^6y^2+4xy^2 ?

OpenStudy (whpalmer4):

24x^2y^6 - 16x^6y^2 + 4xy^2 - (6xy^4*4xy^2) = = 24x^2y^6 - 16x^6y^2 + 4xy^2 - 24x^2y^6 = -16x^6y^2 + 4xy^2 right?

OpenStudy (anonymous):

Yea

OpenStudy (whpalmer4):

okay, so we're doing long division, just not writing it out. we've done the first step, and the first term of our answer is 6xy^4. now we're left with dividing -16x^6y^2 + 4xy^2 by 4xy^2 what is -16x^6y^2 divided by 4xy^2?

OpenStudy (whpalmer4):

-16x^6y^2 ---------- = 4xy^2

OpenStudy (anonymous):

-4x^5y

OpenStudy (whpalmer4):

correct. so multiply -4x^5y by 4xy^2, and subtract that result from what we have left of the original dividend, namely -16x^6y^2 + 4xy^2

OpenStudy (whpalmer4):

that's of course trivial, because we just divided it :-) so our first two terms in our answer are 6xy^4 - 4x^5y now we are left with 4xy^2 , and we need to divide that by 4xy^2.

OpenStudy (anonymous):

-16x^6y^3

OpenStudy (anonymous):

It cancels out

OpenStudy (whpalmer4):

sorry, I said correct, but it wasn't! it should have been -4x^5, not -4x^5y

OpenStudy (whpalmer4):

so our first two terms are 6xy^4 - 4x^5 and we have 4xy^2 divided by 4xy^2 as the part remaining to be done.

OpenStudy (anonymous):

It equals one

OpenStudy (anonymous):

So the answer is 6xy^4 - 4x^5 + 1

OpenStudy (whpalmer4):

right. and we are left with nothing for the remainder, because after we subtract 1*4xy^2 from 4xy^2, the result is 0. let's check our answer: 4xy^2 ( 6xy^4 - 4x^5 + 1) = 24x^2y^6 - 16x^6y^2 + 4xy^2 that's what we want to see!

OpenStudy (whpalmer4):

this is an easy case because we only had 1 term in the divisor (4xy^2). we could have done this one just by eyeballing each term and writing down the answer: 24x^2y^6 - 16x^6y^2 + 4xy^2 -------- -------- ----- = 6xy^4 - 4x^5 + 1 4xy^2 4xy^2 4xy^2

OpenStudy (whpalmer4):

often the divisor will have multiple terms, and in that case, the subtraction is a bit more complicated. but you do the same process.

OpenStudy (anonymous):

If it has multiple terms, can u just add the terms and then divide the other numbers by it ?

OpenStudy (whpalmer4):

for example, x^3 + 3x^2 + 3x + 1 divided by x + 1 x^3/ x = x^2, so x^2 is our first term x^2(x+1) = x^3 + x^2 x^3 + 3x^2 + 3x + 1 - (x^3 + x^2) = 2x^2 + 3x + 1 2x^2 / x = 2x, so 2x is our second term 2x(x+1) = 2x^2 + 2x 2x^2 + 3x + 1 - (2x^2 + 2x) = x + 1 x/x = 1, so 1 is our third term 1(x+1) = x+1 x+1 - (x+1) = 0 so our answer is x^2 + 2x + 1

OpenStudy (anonymous):

Thank you, I understand it now

OpenStudy (whpalmer4):

so when doing this, be careful about - signs in the polynomial you are subtracting. I actually prefer to multiply the one being subtracted by -1 and add instead, I just make fewer mistakes that way.

OpenStudy (anonymous):

Ok thanks